OpenStudy (anonymous):

find an equation of the tangent line to the curve for the given value of t. x=t^2-2t, y=t^2+2t when t=1 find an equation of the tangent line to the curve for the given value of t. x=t^2-2t, y=t^2+2t when t=1 @Mathematics

OpenStudy (anonymous):

ok gimme a sec

OpenStudy (anonymous):

slope of tangent like = 0

OpenStudy (anonymous):

dy/dt = 2t + 2 dx/dt = 2t - 2 dy/dx = (2t + 2)/(2t - 2) plug in 1: (2(1) + 2)/(2(1) - 2) = 4/0 => PROBLEM as dx/dt = 0 at t = 1

OpenStudy (phi):

at t=1, x= 1-2= -1 and y= 1+2 =3. (-1,3) the slope is 4/0 which indicates a vertical line tangent to the point (-1,3). so the equation of the tangent line is x= -1